# Astrophysical signals from the neighboring brane

### v2

@@ Abstract
In this paper we consider the class of extra-dimensional models with two branes and show that each field have to be located on both neighboring branes, thus forming two independent effective fields with different masses. Specific results are illustrated in the framework of a previously developed model. We show that the fermion masses on the two branes are different at the low energies because the fine-tuning of the parameters is necessary only on the observer-populated brane (brane-1). The second brane (brane-2) has no observers and hence the fine-tuning is not obligatory leading to the particle masses of the order of the initial energy scale of the universe formation. The gauge fields are uniformly distributed in the bulk, ensuring charge universality. The Higgs mechanism of masses formation on the both branes is discussed.
Stable, massive charged leptons on brane-2 can be considered as components of the dark matter due to their significantly reduced Compton scattering cross section. Furthermore, we show that inter-brane interaction via photons and gravitons allows the massive fermions on the second brane to be a source of ultra-high energy particles.
@@ 1 Introduction
The idea of multidimensional gravity is a vital tool for obtaining new theoretical results and explaining known phenomena ([Abbott:1984ba, Brown:2013fba, Bronnikov:2009zza, Chaichian:2000az]). Over the course of several decades, numerous significant results have been achieved, and various concepts have been developed. The hierarchy problem is undoubtedly one of the most significant issues (see e.g. [Gogberashvili:1998vx, 1999PhRvL..83.3370R, ArkaniHamed:1998rs]). The problem of the small cosmological constant is discussed in a variety of papers, see e.g. [Krause:2000uj, Bronnikov:2023lej], while the concept of multidimensional inflation is discussed in detail in several sources, including [2002PhRvD..65j5022G, Bronnikov:2009ai, Fabris:2019ecx]. These works assume the extra-dimensional metric is static and stabilization occurs on high-energy scales. The stabilization of extra space as a purely gravitational effect has also been studied in [2002PhRvD..66d4014G, 2003PhRvD..68d4010G, Arbuzov:2021yai].
[…]
@@
For several decades, significant interest has been observed in theories concerning the localization of fields on a three-dimensional hypersurface, known as a brane, within a multidimensional spacetime. Akama was the first who introduced this concept [Akama:1982jy]. The hypothesis of the thick brane was independently formulated by Rubakov and Shaposhnikov [Rubakov:1983bb].
@@
Thin branes, widely discussed in the literature, suffer from some intrinsic problems [Burgess:2015kda]. Therefore, despite hopeful results, they are gradually being replaced by branes with internal structure - thick branes [Bronnikov:2006bu, Chumbes:2011zt, Hashemi_2018, Dzhunushaliev:2019wvv, Bazeia:2022vac, Wan:2020smy]. The latter require special conditions to prove their existence and stability. This could be the choice of a special metric, including the warp factor [Oda_2000] or a special form of the scalar field potential [Bazeia:2022vac]. A purely geometrical approach based on $f(R)$ gravity in 5D is given in [Guo:2023mki], where a considerable number of bibliographic references can be found.
@@
In brane-world scenarios, the key issue is the localization of fields on branes and the recovery of effective 4D gravity. The localization of gauge and spinor fields on branes is a necessary element of any brane model and is a widely discussed topic in the literature. The solutions representing the tensor mode excitations in the presence of a brane are examined in [Cui:2020fiz, Xu:2022xxd] with a focus on the stability problem.
@@
The introduction of the warp factor in a D-dim metric [Oda_2000] generalizes the analysis, especially in the fermion sector, as outlined in [Wan:2023usr], which provides a detailed analysis of this subject in even-dimensional space-time. Fermions in other types of extra metrics are studied in [Gogberashvili_2007, Dantas:2015dca].
@@
The two-brane ideology is often based on the well-known 5D Randall-Sundrum model (RS1), describing two 3D thin branes located at two fixed points of the orbifold. The model postulates a fixed distance between the branes, which is one of its disadvantages. A substantial amount of references can be found in [Olechowski:2024wcf, Tanaka:2000er, Feranie_2010]. The formation of multiple thick branes is considered in, e.g. [Bronnikov:2007kw, Dzhunushaliev:2019wvv]. In this study, we base on the results of our paper [Popov:2024nax] where a new class of branes was discussed. Our analysis there shows that one-brane metrics are exceptions rather than the norm, and we are particularly interested in the set of metrics that describe two-brane configurations. It is shown here that both branes are filled by the same sorts of fields because they have a common origin at the highest energies. In the course of energy reduction, a field is split into two effective fields with different masses depending on the branes properties. Particles located on different branes can interact by exchanging photons. This opens up the possibility of clarifying phenomena such as dark matter [Song:2023xdk] and cosmic rays of the highest energies. Effective superluminal propagation of massless signals in 4D is studied in [Dai:2023zsx].
As was shown in [Bronnikov:2023lej], the Sandwich model of extra dimensions discussed here could serve as the basis for resolving the hierarchy problem, including the issue of the smallness of the cosmological constant.
@@
The paper is structured as follows. In Section 2, we discuss the energy scales suitable for the formation of classical extra-dimensional metrics and the particle content on branes. Section 3 introduces the basis of the Sandwich model of extra dimensions. Section 4 discusses fermions, photons, and the Higgs field distribution over the extra dimensions. We show that the spinor fields split between the two branes. It contains the analysis in which cases the matter located on different branes can be described as independent fields. Section 5 discusses the observable signals from the uninhabited brane and the Higgs mechanism of the mass generation on the both branes. Finally, we give a Conclusion in Section 6.
@@ 2 Spreading of matter on branes in the course of their formation
2.1 Scale of the branes formation
The widely accepted inflationary paradigm implies that the formation of the physical laws takes place above the inflationary scale $H_{I}\sim 10^{14}$eV. Therefore, action describing field dynamics appears at extremely high energies where the quantum fluctuations fade out. The Rundall-Sundrum model with two branes (RS1 model) and its developments [Randall:1999vf] use the second brane as the source of gravity. But there is an aspect which is usually omitted. Indeed, the static two-brane structure does not exist forever, but was formed at high energies. When the universe cools down, two processes occur simultaneously: the extra-dimensional metric describing the branes stabilizes and the fields are randomly settled on them. As a result, the matter fields must be present on all branes.
The two-brane structure is broken by external perturbations if the energy scale is higher than the inverse distance $l$ between the two branes. So we can consider static metrics at the energies where the wavelengths of the 4D fluctuations are much larger than the extra-space size $l_{extra}$. The characteristic wavelengths at the de Sitter stage are of the order of the inverse Hubble parameter $H^{-1}$. Therefore, the estimation $H\equiv H_{D}\ll l_{extra}^{-1}$ provides us with the energy scales where the specific two-brane metric can be considered static.
The D-dim Planck mass $m_{D}$ determines the highest possible energy scale. So the first limit is $l_{extra}^{-1}\ll m_{D}$. Another inequality can be found by using the inflationary concept. It was shown in [Nikulin2019] that a scale of compact extra dimensions cannot be larger than the inflationary scale, $l_{extra}\ll 1/H_{I}$, in order not to destroy the slow rolling during the inflation. Combining these inequalities, we get the interval for the extra-space scale variation
@@ 2.1 Scale of the branes formation
The whole picture looks as follows. The extra-dimensional metric is formed below the energy scale $m_{D}\sim 10^{18}$eV, and is stabilized above the scale $H_{I}\sim 10^{14}$eV. It remains static at lower energies. The initial fluctuations of the fields are responsible for the stochastic distribution of matter on both branes.
@@
The hierarchy of the energy scales and the following fine-tuning are necessary phenomena for the formation of complex structures [2007unmu.book..231D]. The selection of the physical parameters for the fine-tuning on a brane is a rather delicate process. It can be done only for one brane and it is reasonable to choose the brane-1 which is populated by observers. Brane-2 is then populated by all known sorts of particles with masses much heavier than those on brane-1. As discussed above, the physics as a whole and the brane structure in particular are formed at high energy scale in the interval $10^{14}-10^{18}$eV. This means that the natural masses of particles on the brane-2 are of the same order of magnitude.
@@
We come to conclusion that there are all sorts of known particles on the brane-2 which are many order of magnitude more massive compared with those on the brane-1. The matter on the brane-2 is incapable of forming complex structures like stars because of the lack of the fine-tuning.
@@
The potential contribution of such massive particles to the dark matter sector warrants investigation. The following subsection presents a general analysis of this hypothesis, decoupled from specific brane-world configurations. We assume that the gravitational and electromagnetic forces between matter are invariant with respect to the brane on which the matter is located. This assertion, was rigorously established in [Popov:2024nax].
2.2 Dark matter on the uninhabited brane
One of the main goals of the cosmology is to understand the nature of the dark matter. Special efforts have been applied to explain the properties of this phenomenon and the origin of dark matter [Gani:2014lka, 2006PhRvD..73a5001B, 2007PhR...453...29H, 2006PZETF...83..3, Arbuzova:2021etq, Belotsky:2017pcf, Belotsky:2014kca, 2003PhRvD..68j3003B]. Many theoretical models as well as the observational evidences can be found in the reviews [Beylin:2019gtw, Arbey:2021gdg, Misiaszek:2023sxe]. The idea of extra dimensions is also attracted in the framework of this discussion, [Chu:2011be, 2005JCAP...07..001B, 2009MPLA...24..667K, 2006PhRvD..73a5001B]. Some of them are treated the brane idea [Boehmer:2007xh, SHAHIDI_2011] which is natural basis for production of weakly interacted particles. The concept of the mirror world [Beradze:2019yyp, Dubrovich:2021gdn] can be attracted for the same reason.
The most probable content of heavy particles on the brane-2 is the leptons, protons, neutrinos, and their antiparticles. Among these particles, heavy neutrinos emerge as promising candidates for the dark matter component.
Let us return to the two-brane structure discussed above. We have shown that the observers could be on the brane-1, while the heavy electrons with the standard charge are on the brane-2.
A more thorough discussion is necessary to address the implications of charged particles, such as electrons and protons. The central question pertains to the potential role of these charged heavy electrons and protons in the context of dark matter. The primary challenge lies in understanding their interaction with photons, see [Kamada2016EffectsOE]. Such particles, being extremely massive are non-relativistic and could form bonded states like positronium and the hydrogen atom. The positronium life-time $\sim m^{-1}$, ($m$ is the heavy electron mass on the brane-2) and hence decay instantly after its formation. The concentration of the heavy electrons is substantially decreased, yet it remains nonzero. So we have to estimate the Compton effect on brane-2, i.e. the photon scattering on the non-relativistic heavy electron.
The cross section of the Compton effect depends on a particle mass $m$ as
$\sigma\sim\frac{1}{m^{2}}\sim 10^{-26}\text{cm}^{2}.$ (2)
It is easy to see that the heavy electron mass should be really large,
$m_{2}>10^{9}m_{e}\sim 10^{6}\text{GeV},$ (4)
where the value of $\sigma_{e}$ is taken from (2). The mass estimate (4) for heavy electrons on brane-2 allows one to consider them as the dark matter candidate. There are several possible candidates for the role of dark matter: hydrogen, neutrinos, electrons and protons. All of them are located on the brane-2, and hence are very heavy.
The results of this section are based on several assumptions. a) The two-brane metric is the static solution of the classical equations. b) The gravitational and electromagnetic interactions are the same on both branes and uniformly distributed between them. The model elaborated in [Popov:2024nax] is briefly presented below to support points a) and b).
c) Each matter field is necessarily located on two branes and is observed as two independent fields with different masses. This point is discussed in the subsections 4.2 and 4.3.
3 Two-brane metric. Sandwich model of extra dimensions
@@ 3.1 The model choice
The presence of at least two branes is necessary element of this study so that we will base on the model elaborated in [Popov:2024nax]. In that study, we impose the extra-dimensional metric and show that a two-brane structure is formed under some assumptions which are not very restrictive. The metric is supposed in the form
$ds^{2}=\text{e}^{2\gamma(u)}\left[dt^{2}-\text{e}^{2Ht}(dx^{2}+dy^{2}+dz^{2})\right]-du^{2}-r(u)^{2}d\Omega_{n-1}^{2},$ (5)
where $d\Omega_{n-1}^{2}$ is the metric on a unit $n-1$-dimensional sphere. The radial extra coordinate $X^{4}\equiv u$ and $X^{\mu}\equiv x^{\mu},\mu=0,1,2,3.$ The angular coordinate $X^{5}=\theta$ varies in the finite region. Here we consider a 6D space so that $D=4+n,\,n=2$ and the set of extra-dimensional coordinates is $y=\{X^{4},X^{5}\}\equiv\{u,\theta\}$. We do not fix the limits of the radial coordinate $u$ a priori. The Hubble parameter $H$ is arbitrary.
@@
where $g_{\text{D}}\equiv\mathop{\rm det}\nolimits g_{\text{M}\text{N}}$, $\text{M},\text{N}=\overline{1,\text{D}}$, the $n-1$-dimensional manifold $M_{n}$ is assumed to be closed, $f(R)$ is a function of the D-dimensional Ricci scalar $R$, and $m_{\text{D}}$ is the D-dimensional Planck mass.
Remark that there are still no traces of branes, neither in the metric (5) nor in the gravity action (6). Nevertheless, branes are inevitably formed during the cooling of the universe. See [Popov:2024nax]. The two-brane structure reveals as solutions of the classical equations
@@
The static extra-dimensional metrics, being solutions of Eq.7, form an infinite set, as discussed in the papers [Rubin:2014ffa, Rubin:2015pqa, Nikulin2019]. It was shown there that their common form is characterized by two zeros of the metric functions $r(u_{1,2}),\gamma(u_{1,2})$, which is illustrated in Figs.1. The manifold is a compact n-dimensional space of a finite volume.
Figure 1: The typical metric functions, $\gamma(u)$ solid line, $r(u)$ dashed line. The interval is normalized so that the endpoints are $u_{1}=0,\quad u_{2}=\pi$ in the units $m_{D}=1$. Their shapes depends on additional conditions and are characterized by zeros at the endpoints.
@@
The thickness of the brane is related to the fields localization in its vicinity. As was shown in [Bronnikov:2023lej, Popov:2024nax], the classical field distributions along the extra coordinate $u$ are characterized by two areas of localization. One can conclude that the model described above leads to the formation of the two-brane structure and we will refer to it as the Sandwich Model of Extra Dimensions (SMED).

### v3

# Formation of Asymmetrical Two-Brane Structure and its Possible Manifestation

@@ Abstract
In this paper, we consider the class of extra-dimensional models with two branes and show that each field of the Standard Model must be localized on both neighboring branes, whose asymmetry is of great importance. The discussion is conducted in the framework of a previously developed model. Here we show that the Higgs vacuum average is brane-dependent. As the result, fermion masses on the two branes are also different. The second brane (brane-2) lacks observers, eliminating the need for fine-tuning; consequently the particle masses remains of the order of the initial energy scale of the universe formation. Such superheavy charged leptons may serve as a small component of dark matter. Additionally, we show that inter-brane interactions mediated by photons enable massive fermions in brane-2 to act as sources of ultra-high-energy particles. The gauge fields are uniformly distributed in the bulk, ensuring charge universality.
@@
keywords: branes, Higgs, fermion, superheavy particles
  • Natural emergence of the 2-brane structure.
  • The Higgs vacuum average is different across branes.
  • Uninhabited brane generates extremely heavy fermions.
@@ 1 Introduction
The concept of supermassive particles—such as neutrinos and charged leptons—is widely used in modern theoretical physics to explain phenomena like the smallness of neutrino masses [4], the Hubble tension [40], high-energy cosmic rays [8], and dark matter [69], among others. In this article, we explore an extra-dimensional model that inevitably leads to the emergence of supermassive partners of the Standard Model particles.
@@
The multidimensional gravity is a vital tool for obtaining new theoretical results and explaining known phenomena ([1, 25, 28]). Over the course of several decades, numerous significant results have been achieved, and various concepts have been developed. The hierarchy problem is undoubtedly one of the most significant issues (see e.g. [43, 62, 10]). The problem of the small cosmological constant is discussed in a variety of papers, see e.g. [53, 24], while the concept of multidimensional inflation is discussed in detail in several sources, including [44, 38]. These works assume the extra-dimensional metric is static and stabilization occurs on high-energy scales. The stabilization of extra space as a purely gravitational effect has also been studied in [45, 7]. Analytical solutions allow for prompt study of the stability [76, 75, 54] of one-brane configurations in 5D space-time. The stabilization mechanism for the radion in two-brane models was proposed in [31]. The numerical stability of extra dimensions was discussed in our paper [60].
@@
For several decades, significant interest has been observed in theories concerning the localization of fields on a three-dimensional hypersurface, known as a brane, within a multidimensional spacetime. Akama was the first who introduced this concept [3]. The hypothesis of the thick brane was independently formulated by Rubakov and Shaposhnikov [65].
@@
Thin branes, widely discussed in the literature, suffer from some intrinsic problems [26]. Therefore, despite hopeful results, they are gradually being replaced by branes with internal structure - thick branes [22, 30, 47, 37, 13, 72]. The latter require special conditions to prove their existence and stability. This could be the choice of a special metric, including the warp factor [58] or a special form of the scalar field potential [13]. A purely geometrical approach based on $f(R)$ gravity in 5D is given in [46], where a considerable number of bibliographic references can be found.
@@
In brane-world scenarios, the key issue is the localization of fields on branes and the recovery of effective 4D gravity. The localization of gauge and spinor fields on branes is a necessary element of any brane model and is a widely discussed topic in the literature. The solutions representing the tensor mode excitations in the presence of a brane are examined in [32, 75] with a focus on the stability problem.
@@
The introduction of the warp factor in a D-dim metric [58] generalizes the analysis, especially in the fermion sector, as outlined in [73], which provides a detailed analysis of this subject in even-dimensional space-time. Fermions in other types of extra metrics are studied in [42, 34].
@@
The two-brane ideology is often based on the well-known 5D Randall-Sundrum model (RS1), describing two 3D thin branes located at two fixed points of the orbifold. The model postulates a fixed distance between the branes, which is one of its disadvantages. A substantial amount of references and ideas concerning this approach can be found in [59, 71, 39, 33]. The formation of multiple thick branes is considered in, e.g. [21, 37]. In this study, we base on the results of our paper [61] where a new class of branes was discussed. Our analysis there shows that one-brane metrics are exceptions rather than the norm, and we are particularly interested in the set of metrics that describe two-brane configurations.
Here, we present a model based on our latest research that reveals a two-brane structure. Instead of thin branes introduced as a postulate, this model employs thick branes that arise dynamically from a higher-dimensional $f(R)$ gravity background, with matter confined between them.
Moreover, the present scenario predicts double-peaked localization profiles for all massive Standard Model fields— fermions, and the Higgs field. The model generates an intrinsic asymmetry between brane‑1 and brane‑2, resulting in different physical scales in each brane. This asymmetry leads to rich phenomenological implications, such as the existence of super‑heavy partners of Standard Model particles, and naturally gives rise to the concept of a mirror world discussed in, e.g., [17, 36].
In this work, we investigate the consequences of a two-brane configuration in six-dimensional spacetime. We demonstrate that:
- Field Localization and Mass Splitting: Identical Standard Model fields (neutrinos, charged leptons, and quarks) populate both branes. As energy decreases, each fundamental field undergoes spontaneous splitting into two distinct effective fields with brane-dependent mass spectra.
- Higgs Mechanism Asymmetry: The Higgs vacuum expectation value (VEV) differs between the two branes.
- Photons and gravitational waves acting in 4-dim space are distributed uniformly across the extra dimensions (see [61]), enabling interactions between particles on different branes via photon exchange.
- Inter-brane tunneling of particles is possible but highly suppressed.
- Gravitational Interaction: The matter in each brane gravitationally influences the mass distribution of the other brane.
A key result of this study is the existence of superheavy partners of Standard Model particles in brane-2. These partners offer a potential explanation for phenomena such as ultra-high-energy cosmic rays and high-energy neutrinos and constitute a subdominant component of dark matter [41, 69]. At present, we cannot constrain their abundance due to the lack of information about their masses and coupling constants.
@@
The paper is structured as follows. In Section 2, we discuss the energy scales suitable for the formation of classical extra-dimensional metrics and the particle content on branes. Section 3 introduces the basis of our model of extra dimensions. Section 4 discusses fermions, photons, and the Higgs field distribution over the extra dimensions. We show that the spinor fields split between the two branes. It contains the analysis in which cases the matter located on different branes can be described as independent fields and the Higgs mechanism of the mass generation on the both branes. Section 5 discusses the observable signals from the uninhabited brane. Finally, we give a Conclusion in Section 6.
@@ 2 Spreading of matter on branes in the course of their formation
The widely accepted inflationary paradigm implies that the formation of the physical laws takes place above the inflationary scale defined by the Hubble parameter $H_{I}\sim 10^{14}$GeV. Therefore, action describing field dynamics appears at extremely high energies where the quantum fluctuations fade out. The Rundall-Sundrum model with two branes (RS1 model) and its developments [63] use the second brane as the source of gravity. But there is an aspect which is usually omitted. Indeed, the static two-brane structure does not exist forever, but was formed at high energies. When the universe cools down, two processes occur simultaneously: the extra-dimensional metric describing the branes stabilizes and the fields are randomly settled on them. As a result, the matter fields must be present on both branes.
This process is analogous to a second-order phase transition at high temperature. In the non-zero temperature approximation, the effective double-well potential, $V_{\text{eff}}(\Phi,T)$, takes the form
$V_{\text{eff}}(\Phi,T)=V(\Phi)+cT^{2}\Phi^{2}=\left(cT^{2}-\frac{1}{2}\mu^{2}\right)\Phi^{2}+\frac{1}{4}\lambda\Phi^{4}$ (1)
Here, $\mu$ is the mass parameter and $\lambda$ the self-coupling constant of the scalar field $\Phi$. In the high-temperature regime $T\gg\mu$, the effective potential is dominated by the thermal term proportional to $T^{2}\Phi^{2}$, with a positive coefficient $c$ that depends on the couplings of $\Phi$ to other fields in the theory.
In the high-temperature regime, the minimum of $V_{\text{eff}}(\Phi,T)$ is located at $\Phi=0$. As the Universe cools, the coefficient of the $\Phi^{2}$ term changes sign, leading to the formation of two distinct minima. Consequently, the field rolls into one of these minima, creating domains with different vacuum expectation values (VEVs). The random choice of VEV in a given spatial domain results in neighboring regions with distinct VEVs. These domains are separated by a topological wall, which suppresses interactions between field fluctuations in different domains. Specifically, fluctuations are exponentially attenuated when propagating across the wall. The same physical mechanism underlies the process of brane formation.
The two-brane structure is broken by external perturbations if the energy scale is higher than the inverse distance $l_{extra}$ between the two branes. So we can consider static metrics at the energies where the wavelengths of the 4D fluctuations are much larger than the extra-space size $l_{extra}$. The characteristic wavelengths at the de Sitter stage are of the order of the inverse Hubble parameter $H^{-1}$. Therefore, the estimation $H\equiv H_{D}\ll l_{extra}^{-1}$ provides us with the energy scales where the specific two-brane metric can be considered static.
The D-dim Planck mass $m_{D}$ determines the highest possible energy scale. So the first limit is $l_{extra}^{-1}\ll m_{D}$. Another inequality can be found by using the inflationary concept. It was shown in [57] that a scale of compact extra dimensions cannot be larger than the inflationary scale, $l_{extra}\ll 1/H_{I}$, in order not to destroy the slow rolling during the inflation. Combining these inequalities, we get the interval for the extra-space scale variation
@@
The whole picture looks as follows. The extra-dimensional metric is formed below the energy scale $m_{D}\sim 10^{18}$GeV, and is stabilized above the scale $H_{I}\sim 10^{14}$GeV. It remains static at lower energies. The initial fluctuations of the fields are responsible for the stochastic distribution of matter on both branes.
@@
The hierarchy of the energy scales and the following fine-tuning are necessary phenomena for the formation of complex structures [35]. The selection of the physical parameters for the fine-tuning on a brane is a rather delicate process. It can be done only for one brane and it is reasonable to choose the brane-1 which is populated by observers. brane-2 is then populated by all known sorts of particles with masses much heavier than those on brane-1. As discussed above, the physics as a whole and the brane structure in particular are formed at high energy scale in the interval $10^{14}-10^{18}$GeV. This means that the natural masses of particles on the brane-2 are of the same order of magnitude.
@@
We come to conclusion that there are all sorts of known particles on the brane-2 which are many order of magnitude more massive compared with those on the brane-1. The matter on the brane-2 is incapable of forming complex structures because of the lack of the fine-tuning.
@@
The potential contribution of such massive particles to the dark matter sector warrants investigation. The following subsection presents a general analysis of this hypothesis, decoupled from specific brane-world configurations. We assume that the gravitational and electromagnetic forces between matter are invariant with respect to the brane on which the matter is located. This assertion, was rigorously established in [61].
3 Two-brane metric.
@@ 3.1 The model choice
The presence of at least two branes is necessary element of this study so that we will base on the model elaborated in [61]. In that study, we impose the extra-dimensional metric and show that a two-brane structure is formed under some assumptions which are not very restrictive. The metric is supposed in the form
$ds^{2}=g_{AB}dX^{A}dX^{B}=\text{e}^{2\gamma(u)}\left[dt^{2}-\text{e}^{2Ht}(dx^{2}+dy^{2}+dz^{2})\right]-du^{2}-r(u)^{2}d\Omega_{n-1}^{2}.$ (3)
Here we consider a 6D space so that D=4+n,n=2D=4+n,\,n=2 and the set of extra-dimensional coordinates is $y=\{X^{4},X^{5}\}\equiv\{u,\theta\}$. $d\Omega_{n-1}^{2}$ is the metric on a unit n1n-1-dimensional sphere. The angular coordinate $X^{5}=\theta$ varies in the finite region so that the solid angle $d\Omega_{n-1}=d\theta$; $X^{\mu}\equiv x^{\mu},\mu=0,1,2,3$. We do not fix the limits of the radial coordinate $u$ a priori. The Hubble parameter $H$ is arbitrary.
@@
where $g_{\text{D}}\equiv\mathop{\rm det}\nolimits g_{\text{M}\text{N}}$, $\text{M},\text{N}=\overline{1,\text{D}}$, the n1n-1-dimensional manifold $M_{n}$ is assumed to be closed, $f(R)$ is a function of the D-dimensional Ricci scalar $R$, and $m_{\text{D}}$ is the D-dimensional Planck mass.
Remark that there are still no traces of branes, neither in the metric (3) nor in the gravity action (4). Nevertheless, branes are inevitably formed during the cooling of the universe. See [61]. The two-brane structure reveals as solutions of the classical equations
@@
The formation of the two-brane static structure based on $f(R)$ gravity, as in Eq. 5, with the parameters $a=300$ and $c=0.002$, was discussed in our previous papers [66, 57] where additional (boundary) conditions necessary to find a particular solution were also considered. In fact, solutions containing two branes have been found for a wide range of parameters and additional conditions. It was shown there that their common form is characterized by two zeros of the metric functions $r(u_{1,2}),\gamma(u_{1,2})$, which is illustrated in Figs.1. The manifold is a compact n-dimensional space of a finite volume. Here we use a parametric set of metrics with two brane structure
$e^{\gamma(u)}=u^{p}\cdot\sqrt{\sin(u)},$
$r(u)=Au^{q}\cdot\sqrt{\sin(u)}.$ (6)
They retain the above properties being more clear for the analysis. The distance between the branes is the matter of calculations and depends on the additional conditions. We renormalize this distance to the standard interval $(0,\pi)$.
Figure 1: The typical metric functions, $\gamma(u)$ - solid line, $r(u)$ - dashed line. The interval is normalized so that the endpoints are $u_{1}=0,\quad u_{2}=\pi$ in the units $m_{D}=1$. Their shapes depends on additional conditions and are characterized by zeros at the endpoints. The parameter values are $p=1/3,q=1/5,A=2$. As was shown in [61], test particles move toward the endpoints ($u=0$ or $u=\pi$ in this case) along the geodesics derived from the metric containing function $\gamma(u)$.
@@
The thickness of the brane is related to the fields localization in its vicinity. As was shown in [24, 61], the classical field distributions along the extra coordinate $u$ are characterized by two areas of localization. One can conclude that the model described above leads to the formation of the two-brane structure.